OpenStudy (anonymous):

graph the line containing the given pair of points and find the slope (5,2), (-3,-4) m=--?

OpenStudy (anonymous):

m=2+4/(5+3)=3/4

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] pick a point eg (-3,-4) and plug y=(3/4)x -7/4